37.) Show a complete solution to the problem.
If the perimeter of a rectangle is 20 feet and the diagonal is 2 sqrt (13) feet then what are the length and width?

Answer by stanbon(75887)

If the perimeter of a rectangle is 20 feet and the diagonal is 2 sqrt (13) feet then what are the length and width?
------------------------
2(L+W) = 20
L^2 + W^2 = (2sqrt(13))^2
------------------------------
Simplify:
L + W = 10
(10-W)^2 + W^2 = 4*13
------------------------
Solve for W:
100 - 20W + 2W^2 = 52
W^2 - 10W + 24 = 0
(W-4)(W-6) = 0
Width = 4 feet and Length = 6 feet
OR
Width = 6 feet and Length = 4 feet
---------------------
Cheers,
Stan H.